# Fewer supermassive binary black holes in pulsar timing array observations

### v2

@@ Abstract
Observational impact.—When the model closely matches reality, one would expect a reduction of the measurement uncertainty when adding extra data. This is not visible in the original EPTA analysis where the $1\sigma$ range for $(\lg A,\gamma)$ is roughly the same between the 10-year and the 25-year data before our improved noise model. As shown in Figure 1(f), our improved analysis yields a larger measurement uncertainty in the 10-year data, in agreement with our expectation. A decrease in $1\sigma$ uncertainty levels in Figure 1(f) compared to Figure 1(d) indicates that inter-pulsar correlations in the 10-year data provide additional constraints on the background amplitude and spectral index.
@@
The shift of the posteriors in Figure 8 towards larger spectral indices and smaller amplitudes with our improved analysis [^12] The covariance between $\lg A$ and $\gamma$ in Figure 8 is along the line of equal noise power. suggests that the louder pulsar-intrinsic noise with flatter spectra leaks into our measurement of the background strain spectrum when ensemble pulsar noise properties are not modelled.
@@
Because the 10-year data and the 25-year data are not independent data sets, a high degree of consistency is expected. In the original EPTA analysis, maximum-aposteriori $(\lg A,\gamma)$ in the 25-year data differ from those of the 10-year data by around $(0.3,0.8)$. It is visible in red contours across all three panels in Figure 8. A match of the best-fit $(\lg A,\gamma)$ between the 25-year data that does not exhibit Hellings-Downs correlations (Figure 1(b)) and the 10-year data when modelling only temporal correlations and not the Hellings-Downs correlations (Figure 1(f)) is achieved with our improved analysis. However, our improved analysis does not strongly impact best-fit $(\lg A,\gamma)$ in the 10-year data when modelling inter-pulsar correlations. Because $(\lg A,\gamma)$ obtained with only temporal correlations is still expected to match those obtained with temporal and Hellings-Downs correlations, it is also possible that we have not removed all noise model misspecification from the analysis of the EPTA data.
@@
Let us briefly hypothesise about the nature of any other mismodelled noise. We note that the North American Nanohertz Observatory for Gravitational Waves (NANOGrav) has mitigated a tension of the background spectral index with $13/3$ by adopting the Gaussian process model of the dispersion variation noise Nanograv Collaboration (2023a), as in the EPTA analysis EPTA Collaboration and InPTA Collaboration (2023b). Therefore, one potential source of a systematic error could be the mismodelling of the pulsar-specific noise that depends on a radio frequency. A very nearby binary is another example EPTA Collaboration (2023b); Valtolina et al. (2024); Ferranti et al. (2024); Bécsy et al. (2023). Frequency-wise comparison of the inferred strain spectrum against black hole population synthesis models performed earlier by the EPTA (Figure 3 in ref. EPTA Collaboration and InPTA Collaboration (2024)) suggests that the deviation from $\gamma=13/3$ may occur due to excess noise in two frequency bins, $\approx 10^{-8}$ Hz and $\approx 3\times 10^{-8}$ Hz. The rest of the spectrum seems to be consistent with $\gamma=13/3$. The aforementioned potential sources of systematic errors may require better temporal and inter-pulsar correlation models of the data as part of future work.
@@
Methods.–PTAs perform precision measurements of pulse arrival times from millisecond radio pulsars. The likelihood of delays(-advances) $\bm{\delta t}$ for a vector of pulse arrival times $\bm{t}$ is modelled as a multivariate Gaussian distribution $\mathcal{L}(\bm{\delta t}|\bm{\theta})$, where $\bm{\theta}$ is a vector of parameters of models that describe the data. From Bayes’ theorem, it follows that the posterior distribution of model parameters is $\mathcal{P}(\bm{\theta}|\bm{\delta t})=\mathcal{Z}^{-1}\mathcal{L}(\bm{\delta t}|\bm{\theta})\pi(\bm{\theta})$, where $\mathcal{Z}$ is Bayesian evidence, a fully-marginalised likelihood. The term $\pi(\bm{\theta})$ is called a prior probability distribution, a model of how likely it is to find a certain value of $\bm{\theta}$ in Nature. Model selection is performed by computing the ratio of $\mathcal{Z}$ for pairs of models, which is referred to as the Bayes factor. The Bayes factor is equal to the Bayesian odds ratio if both models are assumed to have equal prior odds. A form of the likelihood and the description of the standard analysis methodology can be found in NANOGrav Collaboration (2016). Because the total PTA noise prior is a product of noise priors for every pulsar, PTA data will inform on the distribution of $\bm{\theta}$ in Nature. Our improved analysis introduces hyperparameters $\bm{\Lambda}$ to parametrise priors: $\pi(\bm{\theta}|\bm{\Lambda})\pi(\bm{\Lambda})$. We then perform a numerical marginalisation over $\bm{\Lambda}$. For more details, please refer to the companion paper Goncharov and Sardana (2024).
@@
Data availability.–Second data release of the European Pulsar Timing Array EPTA Collaboration (2023a) is available at zenodo.org and gitlab.in2p3.fr.
@@
Code availability.–The code to marginalise over the uncertainties in pulsar noise priors is available at github.com/bvgoncharov/pta_priors. The PTA likelihood is incorporated in enterprise (Ellis et al., 2020), and posterior sampling is performed using ptmcmcsampler (Ellis and van Haasteren, 2017).
@@
Acknowledgements.–We thank Bruce Allen for helpful comments on the results, contributions to the structuring of this manuscript, and valuable advice on scientific paper writing. We thank Rutger van Haasteren for insightful discussions about hierarchical Bayesian inference. Some of our calculations were carried out using the OzSTAR Australian national facility (high-performance computing) at Swinburne University of Technology. European Pulsar Timing Array is a member of the International Pulsar Timing Array Perera et al. (2019). J. A. acknowledges support from the European Commission (ARGOS-CDS; Grant Agreement number: 101094354). A. C. acknowledges financial support provided under the European Union’s Horizon Europe ERC Starting Grant “A Gamma-ray Infrastructure to Advance Gravitational Wave Astrophysics” (GIGA; Grant Agreement: 101116134).

### v3

@@ III Discussion
When the model closely matches reality, one would expect a reduction of the measurement uncertainty when adding extra data. This is not visible in the original EPTA analysis, where the $1\sigma$ range for $(\lg A,\gamma)$ spans $0.44$ for the 10-year data and $0.39$ for the 25-year data before our improved noise model. As shown in Figure 1(f), our improved analysis yields $0.95$ for the 10-year data and $0.42$ for the 25-year data. A larger reduction in the measurement uncertainty when adding ten extra years of data is in agreement with our expectations. A decrease in $1\sigma$ uncertainty levels in Figure 1(f) compared to Figure 1(d) indicates that inter-pulsar correlations in the 10-year data provide additional constraints on the background amplitude and spectral index. A larger maximum-aposteriori $\lg A$ in Figure 1(d) compared to both Figure 1(b) and Figure 1(f) may also be driven by inter-pulsar correlations.
@@
The shift of the posteriors in Figure 3 towards larger spectral indices and smaller amplitudes with our improved analysis suggests that the louder pulsar-intrinsic noise with flatter spectra leaks into our measurement of the background strain spectrum when the four new sources of noise we point out are not modelled. The covariance between $\lg A$ and $\gamma$ in Figure 3 is along the line of equal noise power. Furthermore, in the early part of the 25-year data, our improved model may ameliorate instrumental noise and a lack of frequency coverage in addition to better modelling of millisecond pulsar spin noise Ferranti et al. (2025a).
@@
Because the 10-year data and the 25-year data are not independent data sets, a high degree of consistency is expected. In the original EPTA analysis, maximum-aposteriori $(\lg A,\gamma)$ in the 25-year data differ from those of the 10-year data by $(0.29,0.76)$. It is visible in red contours across all three panels in Figure 3. A smaller difference between the best-fit $(\lg A,\gamma)$ in the 25-year data and in the 10-year data of $(0.16,0.76)$ is achieved with our improved analysis. When not modelling Hellings-Downs correlations (Figure 1(f)), the difference between the best-fit $(\lg A,\gamma)$ in the 25-year data and in the 10-year data is only $(0.01,0.06)$. Because the best-fit $(\lg A,\gamma)$ obtained with only temporal correlations is still expected to match those obtained with temporal and Hellings-Downs correlations, it is also possible that we have not removed all noise model misspecification from the analysis of the EPTA data.
@@
Let us briefly hypothesise about the nature of any other mismodelled noise. We note that the North American Nanohertz Observatory for Gravitational Waves (NANOGrav) has mitigated a tension of the background spectral index with $13/3$ by adopting the Gaussian process model of the dispersion variation noise Nanograv Collaboration (2023a), as in the EPTA analysis EPTA Collaboration and InPTA Collaboration (2023b). Therefore, one potential source of a systematic error could be the mismodelling of the pulsar-specific noise that depends on a radio frequency. A very nearby binary is another example EPTA Collaboration (2023a); Valtolina et al. (2024); Ferranti et al. (2025b); Bécsy et al. (2023). Frequency-wise comparison of the inferred strain spectrum against black hole population synthesis models performed earlier by the EPTA (Figure 3 in ref. EPTA Collaboration and InPTA Collaboration (2024)) suggests that the deviation from $\gamma=13/3$ may occur due to excess noise in two frequency bins, $\approx 10^{-8}$ Hz and $\approx 3\times 10^{-8}$ Hz. The rest of the spectrum seems to be consistent with $\gamma=13/3$. The aforementioned potential sources of systematic errors may require better temporal and inter-pulsar correlation models of the data as part of future work.
Finally, we calculate evidence for Hellings-Downs correlations in the EPTA data with the revised noise model. Namely, Bayes factor $\mathcal{B}$ in favor of the hypothesis that all pulsars contain Hellings-Downs correlated signal with $(A,\gamma)$ against the hypothesis of the same signal without Hellings-Downs correlations. For 10-year EPTA data, we find $\mathcal{B}=38$ ($\ln\mathcal{B}=3.63$). For the original EPTA noise model, we find a higher value of $\mathcal{B}=59$, consistent with the previously reported value of $60$ in Table 5 in EPTA Collaboration and InPTA Collaboration (2023a). Therefore, our improved model has slightly decreased evidence for Hellings-Downs correlations. It is a combination of the increased measurement uncertainty and, potentially, a fraction of the previously-reported evidence due to the propagation of unmodelled noise into evidence. To evaluate the broad form of inter-pulsar correlations in the 10-year EPTA data, we present Figure 4. It shows the posterior on inter-pulsar correlations, as a function of angular separation between EPTA pulsar pairs, with our revised noise model. The result shows broad consistency with Hellings-Downs correlations [^5] Some of the parameter space in correlation coefficients may be ruled out solely because they do not result in the positive definite covariance matrix of the likelihood. Thus, strictly speaking, flexible correlation models are ill-posed (Rutger van Haasteren, private communication).. For 25-year EPTA data, we find $\mathcal{B}=3$ ($\ln\mathcal{B}=1.17$). For the original EPTA noise model, we find a consistent value ($\mathcal{B}=3$, $\ln\mathcal{B}=0.93$), which is similar to the previously reported value $\mathcal{B}=4$ from Table 5 in EPTA Collaboration and InPTA Collaboration (2023a).
Figure 4: Posterior on inter-pulsar correlations of the common stochastic process in the 10-year EPTA data, assuming our improved noise model, as a function of angular separation between the observed pulsar pairs. Hellings-Downs function of the gravitational wave background is shown as the dashed line.
IV Methods
@@
PTAs perform precision measurements of pulse arrival times from millisecond radio pulsars. The likelihood of delays(-advances) $\bm{\delta t}$ for a vector of pulse arrival times $\bm{t}$ is modelled as a multivariate Gaussian distribution $\mathcal{L}(\bm{\delta t}|\bm{\theta})$, where $\bm{\theta}$ is a vector of parameters of models that describe the data. From Bayes’ theorem, it follows that the posterior distribution of model parameters is $\mathcal{P}(\bm{\theta}|\bm{\delta t})=\mathcal{Z}^{-1}\mathcal{L}(\bm{\delta t}|\bm{\theta})\pi(\bm{\theta})$, where $\mathcal{Z}$ is Bayesian evidence, a fully-marginalised likelihood. The term $\pi(\bm{\theta})$ is called a prior probability distribution, a model of how likely it is to find a certain value of $\bm{\theta}$ in Nature. Model selection is performed by computing the ratio of $\mathcal{Z}$ for pairs of models, which is referred to as the Bayes factor. The Bayes factor is equal to the Bayesian odds ratio if both models are assumed to have equal prior odds. A form of the likelihood and the description of the standard analysis methodology can be found in NANOGrav Collaboration (2016).
Time-correlated noise, which yields stochastic timing delays, can be categorised into achromatic and chromatic. Achromatic noise does not depend on the pulsar’s pulse radio frequency. It is also referred to as spin noise because it represents pulsar rotational irregularities. Chromatic noise (Larsen et al., 2024) is characterised by an amplitude that depends on the radio frequency at which it is observed. Most EPTA pulsars have measurable levels of dispersion measure (DM) noise, and many pulsars are shown to have spin noise (EPTA Collaboration and InPTA Collaboration, 2023b). Epoch-correlated noise, associated with pulse jitter, can influence the inferred strain spectrum of the gravitational wave background at the highest frequencies because of the “white” nature of this noise. Sudden dips with exponential relaxations in timing delays, associated with pulse shape changes (Goncharov et al., 2021b), can also influence parameter estimation for time-correlated signals if not modelled. Noise terms missed in the model, despite evidence, are expected, in many cases, to incorrectly increase the inferred signal amplitude. In particular, the gravitational wave model would attempt to absorb excess noise power.
The priors are our models of how parameters governing pulsar-specific noise are distributed across the pulsars. PTA data provides information about the distribution of $\bm{\theta}$ in Nature, so failing to model this distribution could lead to prior misspecification. To address this, we parametrise prior distributions for amplitudes and spectral indices of pulsar spin and DM noise spectra. Our improved analysis introduces hyperparameters $\bm{\Lambda}$ to parametrise priors: $\pi(\bm{\theta}|\bm{\Lambda})\pi(\bm{\Lambda})$. We then perform a numerical marginalisation over $\bm{\Lambda}$. We use a new procedure of marginalisation over hyperparameters, which is described in Section 2.2.2 of the companion paper Goncharov and Sardana (2025).
Because the gravitational wave background is a stochastic time-correlated signal, it is important to hierarchically model other stochastic time-correlated signals in the data. Pulsar spin noise is described by amplitudes and spectral indices, $(A_{\text{SN}},\gamma_{\text{SN}})$, which are different for every pulsar. Similarly, DM noise is characterised by $(A_{\text{DM}},\gamma_{\text{DM}})$, where the amplitude is referenced to $1400$ MHz. Hierarchical noise model directly impacts posteriors on $(A_{\text{SN}},\gamma_{\text{SN}},A_{\text{DM}},\gamma_{\text{DM}})$. Because the data is only compatible with a certain total amount of timing fluctuations, hierarchical noise inference also impacts posteriors on the gravitational wave background’s amplitude.
V Data availability
@@
Second data release of the European Pulsar Timing Array EPTA Collaboration (2023b) is available at zenodo.org and gitlab.in2p3.fr.
VI Code availability
@@
The code to marginalise over the uncertainties in pulsar noise priors is available at github.com/bvgoncharov/pta_priors. The PTA likelihood is incorporated in enterprise (Ellis et al., 2020), and posterior sampling is performed using ptmcmcsampler (Ellis and van Haasteren, 2017).
VII Acknowledgements
@@
We thank Bruce Allen for helpful comments on the results, contributions to the structuring of this manuscript, and valuable advice on scientific paper writing. We thank Rutger van Haasteren for insightful discussions about hierarchical Bayesian inference. We thank Emily Liepold and anonymous referees for helpful comments on the manuscript. We thank Gabriela Sato-Polito for clarifying the details of Ref. Sato-Polito et al. (2024). This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). Some of our calculations were carried out using the OzSTAR Australian national facility (high-performance computing) at Swinburne University of Technology. European Pulsar Timing Array is a member of the International Pulsar Timing Array Perera et al. (2019). J. A. acknowledges support from the European Commission (ARGOS-CDS; Grant Agreement number: 101094354). A. C. acknowledges financial support provided under the European Union’s Horizon Europe ERC Starting Grant “A Gamma-ray Infrastructure to Advance Gravitational Wave Astrophysics” (GIGA; Grant Agreement: 101116134).